OpenStudy (anonymous):

Andy would like to put a row of decorative tile around his circular pole. the radius of a pole is 5 feet. if the tile cost 12 dollars per foot how much will the tile cost?

OpenStudy (anonymous):

C=2pi*r C=2*3.14*5 C=2*15.7 C=31.4 ft. 31.4*12=$376.8 cost of the tile is this correct ???

OpenStudy (anonymous):

@Mertsj can you check my answer please

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (mertsj):

yes. Good job.

OpenStudy (anonymous):

can you help me on the next one ?

OpenStudy (mertsj):

possibly

OpenStudy (anonymous):

let me know if its blurry

jimthompson5910 (jim_thompson5910):

find the circumference and divide by 4 or find the circumference and multiply by (90/360)

OpenStudy (anonymous):

using 4 as the diameter ?

jimthompson5910 (jim_thompson5910):

4 is the radius

OpenStudy (anonymous):

25.12

OpenStudy (anonymous):

you said divide by 4. so that turns out to be 6.28

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now convert that to a mixed number

jimthompson5910 (jim_thompson5910):

the closest you get is 6 2/7 according to these answer choices

jimthompson5910 (jim_thompson5910):

it's not a perfect fit, but it's close enough

OpenStudy (anonymous):

it says to use pi =22/7 .. does that make a difference ?

jimthompson5910 (jim_thompson5910):

oh yes it makes a slight difference, but you'll still get 6 2/7

OpenStudy (anonymous):

Ok got it ! thanks

jimthompson5910 (jim_thompson5910):

np